# An Autonomous Vision-Based Algorithm for Interplanetary Navigation

### v2

@@ 4.3 Selected Filtering Strategy for the Vision-Based Navigation Algorithm
A non-dimensionalized EKF is selected as the most appropriate filtering approach for the development of a VBN algorithm for CubeSat applications. The selection has been performed in Andreis et al. [9], where the behavior of five different EKFs is analyzed in terms of estimator numerical stability and computational performance. Indeed, it is worth recalling that the autonomous VBN algorithm has to be deployed on a miniaturized processor characterized by limited computation capabilities comparable to the one of a Raspberry Pi [^4] https://www.raspberrypi.com/products/raspberry-pi-4-model-b/, Last Visited on January 2024. The implemented scheme is reported in Table 1, where all the terms are already non-dimensionalized following the approach discussed in Andreis et al. [9].
@@
System State Space
$\dot{\bm{x}}=\bm{f}(\bm{x}(t),t)+\bm{w}$
@@
Propagation Block
$\bm{x}_{p_{k}}=\bm{x}_{c_{k-1}}+\int_{t_{k-1}}^{t_{k}}\bm{f}(\bm{x}(t),t)\textrm{d}t$
$\bm{x}_{c_{0}}=E[\bm{x}_{0}]$
$\bm{P}_{p_{k}}=\bm{P}_{c_{k-1}}+\int_{t_{k-1}}^{t_{k}}{\dot{\bm{P}}\textrm{d}t}$
$\bm{P}_{c_{0}}=E[\bm{x}_{0}\bm{x}_{0}^{\top}]$
Correction Block
$\bm{K}_{k}=\bm{P}_{p_{k}}\bm{H}_{k}^{\top}(\bm{H}_{k}\bm{P}_{p_{k}}\bm{H}_{k}^{\top}+\bm{R}_{k})^{-1}$
@@
Recall that $\bm{x}_{p_{k}}$ is the predicted state vector with error covariance matrix $\bm{P}_{p_{k}}$ at epoch $k$, $\bm{K}_{k}$ the Kalman gain, $\bm{x}_{c_{k}}$ the corrected state vector with error covariance matrix $\bm{P}_{c_{k}}$, $\bm{F}$ the Jacobian of the dynamics model equation, $\bm{h}$ the measurement model equation with Jacobian $\bm{H}_{k}$, $\bm{\nu}_{k}$ the measurement white noise, and $\bm{y}_{k}$ is the external measurement vector $\bm{r}_{\rm{pl}_{k}}$.
The initial probe state vector is perturbed by applying the $3\sigma$ standard deviation rule:
$\bm{x}_{0}=\tilde{\bm{x}}_{0}+3\sqrt{\bm{P}_{0}}\bm{k}$ (51)
where $\tilde{\bm{x}}_{0}$ is the probe nominal state, $\bm{k}$ is a random scalar drawn from the standard distribution. The square root operates on the elements of the initial error covariance matrix $\bm{P}_{0}$, which is defined as:
$\bm{P}_{0}=\rm{diag}(\sigma_{{r}}^{2}I_{3x3},\sigma_{{v}}^{2}I_{3x3},\sigma_{\eta_{\rm{R}}}^{2}I_{3x3},\sigma_{\eta_{\rm{SRP}}}^{2}I_{3x3})$ (52)
where $\sigma_{\eta_{i}}$ is the standard deviation of the respective covariance matrix $\bm{Q}_{\eta_{i}}$. Eventually, two additional procedures are implemented in the navigation filter to face the errors of the IP algorithm:
  1. 1.
    When observations are not acquired due to an IP failure, the state vector and its error covariance matrix are simply propagated until the next step.
  2. 2.
    An innovation-based outlier detection method is applied to reject false positives [46]. In particular, when the absolute value of the innovation term $||\bm{y}_{k}-\bm{h}(\bm{x}_{p_{k}})||$ is greater than $k\sqrt{\bm{M}_{ii}}$ with $\bm{M}=\bm{H}_{k}\bm{P}_{p_{k}}\bm{H}_{k}^{\top}+\bm{R}_{k}$ and $k=3$, the innovation term is set to zero, and the filter correction step is not performed. Indeed, it is preferred to keep an old but good prediction so as not to worsen the estimation.
@@ 5 Results
In this section, primarily, the performance of the IP is tested on a set of randomly generated scenarios. Specifically, a sensitivity analysis is carried on to analyze the IP behavior as the uncertainty on the probe position rises. Following this, the light-time delay analytic equation is verified, and the VBN filter is tested on an interplanetary trajectory. In particular, the impact of light time and light aberration on state estimation is assessed.
@@ 5.1 Image Processing Performance
To validate the IP algorithm performance before adopting it inside the VBN filter, four Monte Carlo (MC) campaigns are carried out to assess the robustness of the pipeline to the variation of the initial spacecraft pose uncertainty. The initial standard deviation of the probe position $\sigma_{\bm{r}}$, which is used to compute the initial filter error covariance matrix (see Eq. 52), is set to $10^{4}$, $10^{5}$, $10^{6}$, and $10^{7}$ km, respectively. In each campaign, the extraction of the beacon location is run for 1000 cases. In each scenario, the $x$ and $y$ components of the spacecraft position are sampled from uniform distribution $U(-3,3)\ \rm{AU}$, while the $z$ component is sampled from $U(-0.07,0.07)\ \rm{AU}$. Note that a translation is then applied to center the spacececraft position in the origin of $\mathcal{I}$ (i.e., the Solar System Barycenter). The $z$-component of the probe position is chosen in a narrower interval as the spacecraft is supposed to lie close to the ecliptic plane. Similarly, the orientation of the probe is determined by sampling Euler angles from a uniform distribution. In particular the right ascension $\alpha$ is sampled from $U(0,2\pi)$, the declination $\delta$ from $U(-0.6,0.6)$, and the twist angle $\phi$ from $U(0,\pi)$. The declination $\delta$ is chosen in a narrower interval as planets are distributed close to the ecliptic plane. The images used to assess the IP performance are generated by adopting an extended and improved version of the sky-field rendering engine illustrated in Bella et al. [47]. In the image simulator, various effects coming from both external and internal sources are modeled and added to deep-space images, such as light effects and detector noises. For what concerns the planet identification step, $\sigma_{\bm{q}_{v}}$ is set equal to 20 arcsec as a result of a statistical analysis conducted on the error obtained in the attitude determination. The planet position uncertainty $\sigma_{\bm{r}_{\rm{pl}}}$ is assumed equal to zero because of the high accuracy with which the planets ephemeris are known. Note that, when other celestial objects are observed to navigate (e.g., asteroids [25] for deep space applications or debris [48] for Earth-based ones), their position and velocity uncertainties may be taken into account in the formulation by using Eq. 48.
to assess the algorithm behavior, the performance indexes adopted for their assessment are the angular error for the attitude determination and the beacon location error for the planet identification step. In detail, the attitude determination error is evaluated by the procedure adopted in Mortari et al. [33]. About the cross-boresight directions, the angular error is determined as
$\sigma_{\textrm{Err}_{\textrm{cb}}}=\cos^{-1}\left({}^{\mathcal{C}}\bm{z}^{\top}\bm{A}_{e}{}^{\mathcal{C}}\bm{z}\right)$ (53)
where $\bm{A}_{e}=\bm{A}_{\textrm{corr}}\bm{A}_{\textrm{true}}^{\top}$ is the error attitude matrix between the estimated and the real attitudes, and ${}^{\mathcal{C}}\bm{z}$ is the boresight direction that corresponds to the body frame third axis. Conversely, the minimum and maximum values of the attitude determination error about the boresight direction are defined as
$\sigma_{{\textrm{Err}_{\textrm{b}}}_{\textrm{min}}}=\cos^{-1}\left(\bm{w}^{\top}\bm{A}_{e}\bm{w}\right)$ (54)
$\sigma_{{\textrm{Err}_{\textrm{b}}}_{\textrm{max}}}=\cos^{-1}\left(\frac{\textrm{trace}(\bm{A}_{e})-1}{2}\right)$ (55)
where
$\bm{w}=\frac{{}^{\mathcal{C}}\bm{z}\times({}^{\mathcal{C}}\bm{z}\times\bm{e})}{||{}^{\mathcal{C}}\bm{z}\times\bm{e}||}$ (56)
with $\bm{e}$ the principal axis of the error attitude matrix $\bm{A}_{e}$.
For what concerns the planet centroid error, this latter is found by evaluating the difference between the position of the centroid identified as planet and the real planet position in the image.
@@
Note that the planet location is detected with a sub-pixel $3\sigma$ accuracy for all the values of $\sigma_{\bm{r}}$. In other terms, the error on the estimated planet location is not dependent on the probe position uncertainty but only on the attitude determination and centroids computation errors. The $3\sigma$ error ellipses in Fig. 2 are obtained from the mean and covariance values reported in Table 2.
Table 2: Mean and covariance of the planet location errors when the probe position uncertainty is known with an accuracy of $10^{4}$, $10^{5}$, $10^{6}$, and $10^{7}$ km
$\sigma_{r}$ [km] $\bm{P}_{\rm{err}}$ $[\rm{px}^{2}]$ $\bm{\mu}_{\rm{err}}$ [px] ${det}(\bm{P})$ $[\textrm{px}^{4}]$
${10^{4}}$ $\begin{bmatrix}0.008&0.0001\\ 0.0001&0.008\\ \end{bmatrix}$ [0.0007, 0.0045]0.00070.0045[-0.0007,\ 0.0045][ - 0.0007 , 0.0045 ] $055.8715e-$ 05
$10^{5}$ $\begin{bmatrix}0.008&0.0001\\ 0.0001&0.008\\ \end{bmatrix}$ [0.0012, 0.0046]0.00120.0046[-0.0012,\ 0.0046][ - 0.0012 , 0.0046 ] $055.8778e-$ 05
$10^{6}$ $\begin{bmatrix}0.008&0\\ 0&0.007\\ \end{bmatrix}$ [0.0011, 0.0052]0.00110.0052[-0.0011,\ 0.0052][ - 0.0011 , 0.0052 ] $055.6399e-$ 05
$10^{7}$ $\begin{bmatrix}0.008&-0.001\\ -0.001&0.007\\ \end{bmatrix}$ [0.0011, 0.0049]0.00110.0049[-0.0011,\ 0.0049][ - 0.0011 , 0.0049 ] $055.4618e-$ 05
@@
The four covariance matrices are characterized by a similar determinant, which is proportional to the area of the ellipse, implying that the accuracy in the estimation of the planet location is not dependent on the uncertainty of the spacecraft position. This feature is one of the key aspects of the adopted pipeline. The results of the IP robustness and attitude determination error are shown in Table 3.
Table 3: Performances of the IP procedure in attitude determination and beacons identification
@@
$\sigma_{r}$ [km] $\sigma_{\textrm{Err}_{\textrm{cb}}}$/$\sigma_{{\textrm{Err}_{\textrm{b}}}_{\textrm{max}}}$ [arcsec] % Wrong Attitude Determination % No Attitude Determination % Correct Beacon Identification
$10^{4}$ 2/13 0.5 0.8 98.75
$10^{5}$ 2/13 0.5 0.6 98.75
$10^{6}$ 2/13 0.5 0.9 98.24
$10^{7}$ 2/14 0.5 1.1 81.28
@@
The algorithm succeeds in determining the probe attitude in over 98% of the scenarios, independently of the probe position uncertainty. The slight variability observed in the no-attitude-determination cases is due to the random behavior of the RANSAC algorithm. The attitude determination is considered wrong when $\sigma_{\rm{Err}_{b_{max}}}>1^{\circ}$. Conversely, the percentage of off-nominal scenarios during planet identification greatly depends on the probe position uncertainty. Indeed, when $\sigma_{\bm{r}}$ increases, the expected planet location is further from the real one, and its covariance ellipse gets larger, leading to a higher probability of planet misidentification. Moreover, the percentage of off-nominal scenarios in planet identification also depends strictly on the success of the attitude determination. Indeed, when attitude determination provides the wrong solution, planet identification fails consequentially. In Table 3, the fifth column represents the percentage of correct planet identification when the attitude determination converges to the right solution. The failure percentage of the beacon identification procedure, when the probe attitude is correctly determined, remains below 2% even with a probe position uncertainty of up to $10^{5}$ km
In conclusion, the precision of the proposed method is independent of the probe position uncertainty, and it relies only on the center-finding performance (i.e., 0.1 pixels). Whereas, the robustness of the IP depends on the attitude determination performance and the probe position uncertainty. Indeed, the planet identification technique precision is not affected by the planet position determination, while its recall is affected by the increase in position uncertainty. It is worth noting that the recall degrades for a high value of the position uncertainties (i.e., 0.01 AU). For completeness sake, Fig. 3 shows some scenarios found during the IP performance assessment where the procedure fails in planet identification.
@@
(b) The attitude is wrongly determined. The planet is mistaken for star 1562
(c) The centroiding algorithm resolves only one centroid, instead of two
(d) Planet wrongly identified. The planet is associated with a wrong spike
@@
In particular, in Figs. 2(a), 2(b), and 2(c) the planet is not found in the image, whereas in Fig. 2(d) the planet is wrongly determined. Regarding Figs. 2(b) and 2(d), the off-nominal cases may be mitigated by applying the RANSAC procedure and the innovation-based outlier detection method, respectively.
@@ 5.2 Filter Results
5.2.1 Assessment of the analytical formulation of the light-time delay
This paragraph specifically investigates the computational time needed to implement the linear model in Eq. 37 and the resulting accuracy in the location of the planet affected by light time. Indeed, it is important to assess the performance of the novel proposed methodology with respect to state-of-the-art techniques in terms of accuracy and computational time. In the analyzed scenario, the light-time effect on the location of the first five planets with respect to an observer at the Solar-System Barycenter in the J200 reference frame is evaluated with the proposed approach and the iterative state-of-the-art solution [45, 19, 9]. The computed locations are then compared to the values found by SPICE [^5] https://naif.jpl.nasa.gov/pub/naif/toolkit_docs/IDL/icy/cspice_spkezr.html, Last Visited on February 2024 which is used as reference. The resulting accuracy in planet location and the computational time required to evaluate the light-time delay with the method proposed in this work are reported in the second and third columns of Table 4. It is worth noting that the computational time is not computed to assess the real performance of the algorithm in flight, but merely for the sake of comparison, given the MATLAB implementation. To fairly compare the performance of the proposed method with the state-of-the-art iterative Newton-Raphson’s approach, the number of iterations required to achieve the same order of accuracy in planet location is evaluated and reported in the sixth column of the table. Furthermore, the computational time needed to compute the light-time delay is outlined in the fifth column. Note that computational times have been computed by taking the average values over 1000 runs in both methods. By comparing the computational times, the proposed approach performs always better than the iterative one. Consequently, with equivalent accuracy, the proposed approximation proves to be significantly more computationally efficient than the iterative method. For the sake of completeness, the computational time needed to achieve a less stringent accuracy level of 0.04 pixels (i.e., one order below the accuracy of the IP) is also computed for the Mars case. The iterative method still takes 5 ms with 18 iterations, further supporting the preference for the proposed approach.
Table 4: Obtained accuracy in the estimation of the planet position affected by light time and required computational time to compute the time delay with the proposed method and the iterative approach
Proposed method Iterative approach
Accuracy [px]
CPU time [ms]
Accuracy [px]
CPU time [ms]
N iterations [-]
Mercury 2.32e-05 0.019 2.32e-05 3 18
Venus 8.15e-07 0.1 8.18e-07 48 23
Earth 4.37e-07 0.086 4.38e-07 53 24
Mars 1.28e-06 0.025 1.28e-06 6 21
Jupiter 5.42e-05 0.088 5.19e-05 37 22
@@ 5.2.2 Navigation Concept of Operations
In the study case, a CubeSat estimates its position and velocity by tracking visible planets over an interplanetary transfer. The spacecraft alternates observation windows, where an asynchronous tracking of the optimal pair of planets is performed, to only-propagation windows, where the filter only propagates the probe state as no external observations are acquired. The navigation Concept of Operations (CONOPS) is shown in Fig. 4. The probe tracks the first planet of the optimal pair, which is selected at the beginning of the navigation cycle, then performs a slew maneuver to point to the second planet, during which no observations are acquired, and it observes this later. Eventually, the estimation is propagated until the following observation window starts.
@@
At every time step of the planet observation, an image is generated using an improved version of the deep-space sky simulator in Bella et al. [47]. The simulator models the effects caused by the lights (i.e., light-time and light-aberration effects) on the centroids positions and by the impact of cosmic rays hitting the sensor frame. The sky simulator renders the image by taking into input the true probe pose and velocity. Since the attitude control system is not simulated in this work, the true probe orientation is computed by evaluating the desired pointing direction needed to acquire the planet at the center of the image and adding a random perturbation to it, which simulates the spacecraft jitter effect and the attitude knowledge error. Since the probe position is known with a given uncertainty (up to $10^{5}$ km in this work), the beacon location will be not perfectly centered in the image, but still be contained in it, which is a sufficient condition to let the IP pipeline extract the planet observation.

### v3

@@ 4.3 Selected Filtering Strategy for the Vision-Based Navigation Algorithm
A non-dimensionalized EKF is selected as the most appropriate filtering approach for the development of a VBN algorithm for CubeSat applications. The selection has been performed in [9], where the behavior of five different EKFs has been analyzed in terms of estimator numerical stability and computational performance. Indeed, it is important to remark that the autonomous VBN algorithm has to be deployed on a miniaturized processor characterized by limited computation capabilities comparable to the one of a Raspberry Pi. The implemented scheme is reported in Table 1, where all the terms are already non-dimensionalized following the approach discussed in [9].
@@
System State Space $\dot{\bm{x}}=\bm{f}(\bm{x}(t),t)+\bm{w}$
@@
Propagation Block $\bm{x}_{p_{k}}=\bm{x}_{c_{k-1}}+\int_{t_{k-1}}^{t_{k}}\bm{f}(\bm{x}(t),t)\textrm{d}t$ $\bm{x}_{c_{0}}=E[\bm{x}_{0}]$
$\bm{P}_{p_{k}}=\bm{P}_{c_{k-1}}+\int_{t_{k-1}}^{t_{k}}{\dot{\bm{P}}\textrm{d}t}$ $\bm{P}_{c_{0}}=E[\bm{x}_{0}\bm{x}_{0}^{\top}]$
Correction Block $\bm{K}_{k}=\bm{P}_{p_{k}}\bm{H}_{k}^{\top}(\bm{H}_{k}\bm{P}_{p_{k}}\bm{H}_{k}^{\top}+\bm{R}_{k})^{-1}$
@@
Here, $\bm{x}_{p_{k}}$ is the predicted state vector with error covariance matrix $\bm{P}_{p_{k}}$ at epoch $k$, $\bm{K}_{k}$ the Kalman gain, $\bm{x}_{c_{k}}$ the corrected state vector with error covariance matrix $\bm{P}_{c_{k}}$, $\bm{F}$ the Jacobian of the dynamics model equation, $\bm{h}$ the measurement model equation with Jacobian $\bm{H}_{k}$, $\bm{\nu}_{k}$ the measurement white noise, and $\bm{y}_{k}$ the external measurement vector.
Moreover, two additional procedures are implemented in the navigation filter to face the errors of the IP algorithm: 1) When observations are not acquired due to an IP failure, the state vector and its error covariance matrix are simply propagated until the next step; 2) an innovation-based outlier detection method is applied to reject false positives [41]. In particular, when the absolute value of the innovation term ($||\bm{r}_{\rm{pl}_{k}}-\bm{h}(\bm{x}_{p_{k}})||$) is greater than $k\sqrt{\bm{M}_{ii}}$ with $\bm{M}=\bm{H}_{k}\bm{P}_{p_{k}}\bm{H}_{k}^{\top}+\bm{R}_{k}$ and $k=3$, the innovation term is set to zero, and the filter correction step is not performed. Indeed, it is preferred to keep an old but good prediction so as not to worsen the estimation.
@@ 5.1 Image Processing Performance
To validate the IP algorithm performance before adopting it inside the VBN filter, four Monte Carlo campaigns are carried out where the initial uncertainty of the probe position $\sigma_{\bm{r}}$ is set to $10^{4}$, $10^{5}$, $10^{6}$, and $10^{7}$ km, respectively. In each campaign, the extraction of the beacon position projection is run for 1031 scenarios, wherein at least one planet is present, out of the 3000 analyzed. In each scenario, the position of the spacecraft is selected by randomly sampling a Gaussian distribution with $\sigma_{x}=\sigma_{y}=3\rm{AU}$ and $\sigma_{z}=0.07\rm{AU}$ and centered in the origin of $\mathcal{N}$. The $z$-component of the probe position is chosen in a narrower interval as the spacecraft is supposed to lie close to the ecliptic plane. Similarly, the orientation of the probe is assigned by randomly sampling a normal distribution $\alpha$, $\delta$, and $\phi$ in the $3\sigma$ intervals $[0,2\pi]$, $[-0.6,0.6]$, and $[0,2\pi]$, respectively. The declination $\delta$ is chosen in a narrower interval as planets are distributed close to the ecliptic plane. For what concerns the planet detection step, $\sigma_{\bm{q}_{v}}$ is set equal to 20 arcsec as a result of a statistical analysis conducted on the error obtained in the attitude determination. The planet position uncertainty $\sigma_{\bm{r}_{\rm{pl}}}$ is assumed equal to zero because of the high accuracy with which the planets ephemeris are known.
@@
Note that the planet position projection is detected with a sub-pixel $3\sigma$ accuracy for all the values of $\sigma_{\bm{r}}$. In other terms, the error on the estimated planet position projection is not dependent on the probe position uncertainty but only on the attitude determination and centroids computation errors. When the probe position uncertainty increases, the scenarios where the beacon projection error is over 0.3 px seem to be filtered out. Indeed, in these cases, the IP algorithm may select a different wrong spike as the expected position projection becomes far from the real one. As a result, the error norm becomes greater than the threshold value set for the assessment of the IP performance. This case is considered to be a failure of the IP pipeline and it is not represented in the PDF. Note that this condition is only applied during the IP testing, therefore, it is not present when the filter is run. The $3\sigma$ error ellipses in Fig. 2 are obtained from the mean and covariance values reported in Table 2.
Table 2: Mean and covariance of the planet position projection errors for the four values of $\sigma_{\bm{r}}$
$\sigma_{\bm{r}}$ [km] $10^{4}$ $10^{5}$ $10^{6}$ $10^{7}$
$\bm{P}_{\rm{err}}$ $[\rm{px}^{2}]$ $\begin{bmatrix}0.008&0.001\\ 0.001&0.007\\ \end{bmatrix}$ $\begin{bmatrix}0.011&0.002\\ 0.002&0.007\\ \end{bmatrix}$ $\begin{bmatrix}0.013&0.006\\ 0.006&0.011\\ \end{bmatrix}$ $\begin{bmatrix}0.007&0.001\\ 0.001&0.005\\ \end{bmatrix}$
$\bm{\mu}_{\rm{err}}$ [px] [0.0014;-0.0003] [0.0001;-0.0034] [-0.0005; -0.0045] [-0.0005;-0.0041]
$\text{det}(\bm{P})$[px4] 5.9e-05 7.05e-05 9.94e-05 7.05e-05
@@
The four covariance matrices are characterized by a similar determinant, which is proportional to the area of the ellipse, implying that the precision in the planet determination is not dependent on the uncertainty of the spacecraft position. This feature is one of the advantages of the proposed pipeline for planet detection in deep-space images. The results of the IP robustness and attitude determination error are shown in Table 3 for the 962 cases in which a correct or wrong attitude solution is found.
Table 3: Algorithm Performance
@@
$\sigma_{r}$ [km] $\sigma_{\rm{ErrRot}}$ [arcsec] % Wrong Beacon Detection (of 962 cases) % Wrong Beacon Detection with Right Attitude Determination (of 962 cases)
$10^{4}$ 14.77 4.57 (44 cases) 0.42 (4 cases)
$10^{5}$ 15.18 4.68 (45 cases) 0.52 (5 cases)
$10^{6}$ 15.21 7.69 (74 cases) 3.33 (32 cases)
$10^{7}$ 15.48 29.83 (287 cases) 25.47 (245 cases)
@@
The percentage of off-nominal scenarios during planet identification greatly depends on the probe position uncertainty. Indeed, when $\sigma_{\bm{r}}$ increases, the expected planet position projection is further from the real position projection, and its uncertainty ellipse is bigger, which leads to a higher probability of planet misidentification. Moreover, the percentage of off-nominal scenarios in planet detection also depends strictly on the success of the attitude determination. Indeed, when attitude determination provides the wrong solution, planet detection fails consequentially. In Table 3, the third column represents the total number of cases of wrong planet identification when the attitude determination converges to a right or wrong solution. Instead, the last column represents the number of cases of wrong planet identification when the attitude determination converges to the correct solution. The failure percentage of the beacon detection procedure when the probe attitude is correctly determined is lower than 1% with a probe position uncertainty up to $10^{5}$ km.
@@
(b) The attitude is wrongly determined. The planet is mistaken for star 1562.
(c) The centroiding algorithm resolves only one centroid, instead of two.
(d) Planet wrongly identified. The planet is associated with a wrong spike.
Figure 3: Scenarios in which the IP pipeline fails in the planet detection.
@@ 5.2.1 Navigation Concept of Operations
In the study case, a CubeSat estimates its position and velocity by tracking visible planets over an interplanetary transfer. The spacecraft alternates observation windows, where an asynchronous tracking of the optimal pair of planets is performed, to only-propagation windows, where the filter only propagates the probe state as no external observations are acquired. The navigation CONOPS is shown in Fig. 4. The probe tracks the first planet of the optimal pair, which is selected at the beginning of the navigation cycle, then performs a slew maneuver to point to the second planet, during which no observations are acquired, and it observes this later. Eventually, the estimation is propagated until the following observation window starts.
@@
At every time step of the planet observation, an image is generated using an improved version of the deep-space sky simulator in [43]. The simulator models the effects caused by the lights, i.e., light-time and light-aberration effects, on the centroids’ positions and by the impact of cosmic rays hitting the sensor frame. The sky simulator renders the image by taking into input the true probe pose and velocity. Since the attitude control system is not simulated in this work, the true probe orientation is computed by evaluating the desired pointing direction needed to acquire the planet at the center of the image and adding a random perturbation to it, which simulates the spacecraft jitter effect and the attitude knowledge error. Since the probe position is known with a given uncertainty (up to $10^{5}$ km in this work), the beacon projection will be not perfectly centered in the image, but still contained in it, which is a sufficient condition to let the IP pipeline extract the planet observation.